Grade 12Mechanics

Two racing cars of masses m1 and m2 are moving in circles of raddi r1 and r2 respectively. Their speeds are such that each makes a complete circle of radii r1 and r2 respectively.Their speeds are such that each makes a complete circle in time t. the ratio of the angular speeds of the first car to the second is
a)1:1 b)m1:m2
c)r1:r2 d)m1m2:r1r2

prateek
8 Years agoGrade 12

1 Answer

sourav
ApprovedApproved Tutor Answer8 Years ago
Since t1,= t2 = t therefore we can consider,
\frac{\omega_1}{\omega _2} = \frac{\frac{2 \pi}{t}}{\frac{2 \pi}{t}} =1:1.